OpenStudy (anonymous):

Find the slope of the given line. 7x - 2y = 4 m=?

OpenStudy (anonymous):

-2y=-7x+4 y=7/2 x -2 slope m= 7/2 (coefficient in a front of x)

OpenStudy (anonymous):

thank you!